OpenStudy (anonymous):

rationalize [sqrt(4+x^5)-2]/(x^5) and show me the steps. Whenever I rationalize the numerator it goes to 0, which is bad since I'm trying to find the limit of this function.

OpenStudy (dumbcow):

so you multiply by sqrt(4+x^5) +2 right

OpenStudy (anonymous):

\[\frac{\sqrt{4+x^5}-2}{x^5}\]like that ?

OpenStudy (anonymous):

yes satellite

OpenStudy (anonymous):

i think you should get a 1 up top, not a zero

OpenStudy (dumbcow):

you should get --> (4+x^5) - 4 = x^5

OpenStudy (anonymous):

riight because the x^5 s cancel to 1, not 0.

OpenStudy (anonymous):

Do'oh!

OpenStudy (anonymous):

lol how'd i guess?

OpenStudy (dumbcow):

haha :)

OpenStudy (anonymous):

thanks guys.